Analyzing, Forecasting, and Interpreting Income Statement and Balance Sheet
Following are the income statement and balance sheet of Whole Foods Market, Inc.

Income Statement,
For Years Ended (in $ 000s)
2010 2009
Sales $ 9,005,794 $ 8,031,620
Cost of goods sold and occupancy costs 5,870,393 5,277,310
Gross profit 3,135,401 2,754,310
Direct store expenses 2,375,716 2,145,809
General administrative expenses 272,449 243,749
Pre-opening expenses 38,044 49,218
Relocation, store closures and lease termination costs 11,217 31,185
Operating income 437,975 284,349
Interest expense (33,048) (36,856)
Investment and other income 6,854 3,449
Income before income taxes 411,781 250,942
Provision for income taxes 165,948 104,138
Net income $ 245,833 $ 146,804
Balance Sheet,($ 000s) 2010 2009
Assets
Cash and cash equivalents $ 131,996 $ 430,130
Short-term investments-available-for-sale securities 329,738 --
Restricted cash 86,802 71,023
Accounts receivable 133,346 104,731
Merchandise inventories 323,487 310,602
Prepaid expenses and other current assets 54,686 51,137
Deferred income taxes 101,464 87,757
Total current assets 1,161,519 1,055,380
Property and equipment, net of accum depr and amortization 1,886,130 1,897,853
Long-term investments--available-for-sale securities 96,146 --
Goodwill 665,224 658,254
Intangible assets, net of accumulated amortization 69,064 73,035
Deferred income taxes 99,156 91,000
Other assets 9,301 7,866
Total assets $ 3,986,540 $ 3,783,388
Liabilities and shareholders' equity
Current installments of long-term debt and capital lease obligations $ 410 $ 389
Accounts payable 213,212 189,597
Accrued payroll, bonus and other benefits due team members 244,427 207,983
Dividends payable -- 8,217
Other current liabilities 289,823 277,838
Total current liabilities 747,872 684,024
Long-term debt and capital lease obligations, less current installments 508,288 738,848
Deferred lease liabilities 294,291 250,326
Other long-term liabilities 62,831 69,262
Total liabilities 1,613,282 1,742,460
Series A redeemable preferred stock; $0.01 par value -- 413,052
Shareholders' equity:
Common stock, no par value 1,773,897 1,283,028
Accumulated other comprehensive income 791 (13,367)
Retained earnings 598,570 358,215
Total shareholders' equity 2,373,258 1,627,876
Total liabilities and shareholders' equity $ 3,986,540 $ 3,783,388

Forecast Whole Foods Market's 2011 income statement using the following relations (assume "no change" for accounts not listed). ($ 000s)

Sales growth 10.0%
Cost of good sold and occupancy costs/Sales 65.2%
Direct store expenses/Sales 26.4%
General and administrative expenses/Sales 3.0%
Pre-opening expenses/Sales 0.4%
Relocation, store closures and lease termination costs $--
Interest expense $33,048
Investment and other income $6,854
Provision for income taxes/Pretax income 40.3%
  • Round answers to the nearest whole number.

  • Do not use negative signs with your answers.

Whole Foods Market, Inc.
Income Statement, For Years Ended ($ 000s)
2011 Estimated
Sales Answer
Cost of goods sold and occupancy costs Answer
Gross profit Answer
Direct store expenses Answer
General and administrative expenses Answer
Pre-opening expenses Answer
Relocation, store closures and lease termination costs Answer
Operating income Answer
Interest expense Answer
Investment and other income Answer
Income before income taxes Answer
Provision for income taxes Answer
Net income Answer

Forecast Whole Foods Market's 2014 balance sheet using the following relations (assume "no change" for accounts not listed). ($ 000s)

Restricted cash $86,802
Short-term investments-available-for-sale securities $329,738
Accounts receivable/Sales 1.5%
Merchandise inventories/Sales 3.6%
Prepaid expenses and other current assets/Sales 0.6%
Deferred income taxes (current) $101,464
Long-term investments--available-for-sales securities $96,146
Goodwill $665,224
Intangible assets, net of accumulated amortization $69,064
Deferred income taxes (noncurrent) $99,156
Other Assets/Sales 0.1%
Current installments of long-term debt and capital lease obligations $410
Accounts payable/Sales 2.4%
Accrued payroll, bonus and other benefits due team members/Sales 2.7%
Other current liabilities/Sales 3.2%
Deferred lease liabilities $294,291
Other long-term liabilities/Sales 0.7%
Series A redeemable preferred stock $--
Common stock $1,773,897
Accumulated other comprehensive income $791
Capital expenditures/Prior year net PPE 13.5%
Depreciation & amortization/Prior year net PPE 14.5%
Dividends/Net income 3.5%
  • Round answers to the nearest whole number.

Whole Foods Market, Inc.
Balance Sheet, ($000s)
2011 Estimated
Assets
Cash and cash equivalents Answer
Short-term investments-available-for-sale securities Answer
Restricted cash Answer
Accounts receivable Answer
Merchandise inventories Answer
Prepaid expenses and other current assets Answer
Deferred income taxes Answer
Total current assets Answer
Property and equipment, net of accum depr and amortization Answer
Long-term investments--available-for-sale securities Answer
Goodwill Answer
Intangible assets, net of accumulated amortization Answer
Deferred income taxes Answer
Other assets Answer
Total assets Answer
Liabilities and shareholders' equity
Current installments of long-term debt and capital lease obligations Answer
Accounts payable Answer
Accrued payroll, bonus and other benefits due team members Answer
Dividends payable Answer
Other current liabilities Answer
Total current liabilities Answer
Long-term debt and capital lease obligations, less current installments Answer
Deferred lease liabilities Answer
Other long-term liabilities Answer
Total liabilities Answer
Series A redeemable preferred stock, $0.01 par value Answer
Shareholders' equity:
Common stock, no par value Answer
Accumulated other comprehensive income Answer
Retained earnings Answer
Total shareholders' equity Answer
Total liabilities and shareholders' equity Answer
Answer #1

Answer:

Income Statement for Year ended 2011 E Formulas for reference
Sales 9,906,373 =9005794*(1+10%)
Cost of goods sold and occupancy costs 6,458,955 =9906373*65.2%
Gross Profit 3,447,418 =9906373-6458955
Direct store exp 2,615,282 =9906373*26.4%
G&A Exp 297,191 =9906373*3%
Pre-opening exp 39,625 =9906373*0.4%
relocation store closures and leas termination costs 0
Operating Income 495,319 =3447418-(2615282+297191+39625)
Interest Exp 33,048
Investment and other income 6,854
Income before income tax 469,125 =495319-33048+6854
Provision for income taxes 189,057 =469125*40.3%
Net Income 280,067 =469125-189057
Whole foods Market Inc
Balance Sheet
Assets 2011 E
Cash and cash Balance 459,013
Short term Investment available for sale securitites 329,738
Restricted Cash 86,802
Accounts Receivable 148,596 =9906373*1.5%
Merchandise Inventories 356,629 =9906373*3.6%
Prepaid Exp and Other Current Assets 59,438 =9906373*0.6%
Deferred Income Taxes 101,464
Total Current Assets 1,541,680
Property And Equipment Net of Dept 1,612,641 =1886130*(1-14.5%)
Long Term Investment Available for Sale Securities 96,146
Goodwill 665,224
Intangible Assets of Acc Amount 69,064
Deferred Income Taxes non current 99,156
Other Assets 9,906 =9906373*0.1%
Total Assets 4,093,818
Liability and Shareholders Equity
Current Installment of long term debt and capital lease obligation 410
Accounts payable 237,753 =9906373*2.4%
Accured payroll bonus and other benefit 267,472 =9906373*2.7%
Dividends payable 9,802 =280067*3.5%
Other current liabilities 317,004 =9906373*3.2%
Total Current Liabilities 832,441
Long term debt capital lease obligation less current installment 254,218 =(1886130*13.5%)-410
Deferred lease liability 294,291
Other long term liabilities 69,345 =9906373*0.7%
Total Liabilities 1,450,295
Series a redemable P Stocks
Shareholders Equity
Common Stock No par value 1,773,897
Accumulated Other Comprehnsive Income 791
Retained Earnings 868,835
Total Shareholders Equity 26,43,523
Total Liabilities and Shareholders Equity 4,093,818